# The Angular Momentum Penrose Inequality

### v3

@@ 4.3 Axisymmetric Setting
At the poles, $\rho_{\Sigma}=0$, so $\mathcal{T}=O(s\cdot 0)=0$. The explicit constant follows from $\rho_{\Sigma}\leq\rho_{\max}$. ∎
@@
Lemma 4.16 (Perturbation Stability for Blow-Up Asymptotics).
@@
$\|\Phi(v_{1})-\Phi(v_{2})\|_{W^{2,2}_{\beta}}$ $\leq C_{L}\bigl(\|N[v_{1}]-N[v_{2}]\|_{W^{0,2}_{\beta}}+\|\mathcal{T}[f_{0}+v_{1}]-\mathcal{T}[f_{0}+v_{2}]\|_{W^{0,2}_{\beta}}\bigr).$ (32)
@@
Remark 4.17 (Explicit Constant Dependencies).
The perturbation stability argument involves the following explicit constants, with quantitative formulas in terms of the spectral data:
  • $C_{L}=\|L_{0}^{-1}\|_{W^{0,2}_{\beta}\to W^{2,2}_{\beta}}$: The Fredholm inverse bound admits the explicit estimate
    $C_{L}\leq\frac{C_{\mathrm{elliptic}}}{(\gamma_{0}-|\beta|)(\gamma_{0}+|\beta|)}=\frac{C_{\mathrm{elliptic}}}{\gamma_{0}^{2}-\beta^{2}},$ (34)
    where $\gamma_{0}=\sqrt{\lambda_{1}(L_{\Sigma})/8}$ is the smallest positive indicial root determined by the principal eigenvalue $\lambda_{1}(L_{\Sigma})>0$ of the MOTS stability operator, and $C_{\mathrm{elliptic}}$ is a universal constant from standard elliptic theory depending only on the dimension and ellipticity constants. For $\beta$ chosen as $\beta=-\gamma_{0}/2$, we obtain:
    $C_{L}\leq\frac{4C_{\mathrm{elliptic}}}{3\gamma_{0}^{2}}=\frac{32C_{\mathrm{elliptic}}}{3\lambda_{1}(L_{\Sigma})}.$
    This shows $C_{L}$ is inversely proportional to the spectral gap $\lambda_{1}(L_{\Sigma})$: more stable MOTS yield smaller $C_{L}$ and better perturbation control.
  • $C_{N}\leq C\|\nabla^{2}f_{0}\|_{L^{\infty}_{\text{loc}}}$: bounded by the $C^{2}$ norm of the unperturbed solution. By the Han–Khuri blow-up analysis [29, Proposition 4.5], $\|\nabla^{2}f_{0}\|_{L^{\infty}(K)}\leq C(K,|\theta^{-}|)$ on any compact set $K\subset M\setminus\Sigma$, where $|\theta^{-}|=2C_{0}$ is the inner null expansion magnitude.
  • $C_{\mathcal{T}}=C_{\omega,\infty}\cdot\rho_{\max}^{2}/C_{0}$: bounded by the twist 1-form norm $C_{\omega,\infty}=\sup_{\mathcal{Q}}|\omega|$, maximum orbit radius $\rho_{\max}=\sup_{\Sigma}\rho$, and the blow-up coefficient $C_{0}=|\theta^{-}|/2>0$.
  • $\delta=\min\left(\frac{1}{4C_{L}C_{N}},\sqrt{\frac{1}{2C_{L}C_{\mathcal{T}}}}\right)$: the ball radius for the contraction map. Substituting the explicit bounds:
    $\delta\geq\min\left(\frac{3\lambda_{1}(L_{\Sigma})}{128C_{\mathrm{elliptic}}C_{N}},\sqrt{\frac{3\lambda_{1}(L_{\Sigma})}{64C_{\mathrm{elliptic}}C_{\mathcal{T}}}}\right).$
For axisymmetric vacuum data with strictly stable MOTS ($\lambda_{1}(L_{\Sigma})>0$), all these constants are finite and explicitly computable from the initial data $(M,g,K)$. The formulas show that the perturbation argument becomes quantitatively stronger (larger $\delta$) when: (i) the MOTS is more stable (larger $\lambda_{1}$), (ii) the twist is weaker (smaller $C_{\omega,\infty}$), and (iii) the horizon is farther from extremality (smaller $\rho_{\max}$).
@@
$L_{\mathrm{axi}}v=\frac{1}{\sqrt{1+|\nabla f|^{2}}}\left[\Delta v-\frac{\nabla^{i}f\nabla^{j}f}{1+|\nabla f|^{2}}\nabla_{ij}v-(|A_{\Gamma}|^{2}+\mathrm{Ric}(\nu_{\Gamma},\nu_{\Gamma}))v\right]+\text{(K-terms)}+D\mathcal{T}|_{f}\cdot v,$
@@
$0$ $=\int_{\mathcal{C}_{T}}v\cdot L_{\mathrm{axi}}v\,dV_{\bar{g}}$ (37)
$=\int_{\mathcal{C}_{T}}\left[-|\nabla v|^{2}+\mathcal{V}v^{2}+O(e^{-\beta_{0}t})|v|^{2}+O(e^{-t})|v||\nabla v|\right]dV_{\bar{g}}+\text{(boundary terms)}.$ (38)
@@
  • The blow-up asymptotics $|\nabla f|\sim C_{0}/s$ (from Han–Khuri [29, Prop. 4.5]);
@@
  • The $\rho^{2}$ scaling of the twist term: $\mathcal{T}\propto\rho^{2}$, which vanishes at the poles where $\rho=0$ (Lemmas 4.6 and 4.8).
@@ 5.1 The Conformal Equation
The Jang manifold $(\bar{M},\bar{g})$ has two distinct asymptotic regions requiring different function space frameworks:
  1. (i)
    Asymptotically flat end: Weighted Hölder spaces $C^{k,\beta}_{-\tau}$ with polynomial weight $r^{-\tau}$ (Definition 4.1);
  2. (ii)
    Cylindrical end: Weighted Sobolev spaces $W^{k,p}_{\beta}$ with exponential weight $e^{\beta t}$ (Definition 5.1).
These frameworks are compatible on the transition region $\{R_{0}\leq r\leq 2R_{0}\}$ (equivalently $\{0\leq t\leq T_{0}\}$) in the following sense: by Sobolev embedding, $W^{k+1,2}_{\beta}\hookrightarrow C^{k,\beta}$ locally, and both norms are equivalent (up to constants depending on $R_{0}$) on the compact overlap region. This allows elliptic estimates to be “glued” across the transition using standard partition-of-unity arguments. The key point is that the Fredholm index is determined by the asymptotic behavior at both ends, not the transition region.
@@
The angular-momentum-modified Lichnerowicz equation is:
@@
We verify the sign conventions in the AM-Lichnerowicz equation:
  1. (i)
    Conformal transformation formula: Under $\tilde{g}=\phi^{4}\bar{g}$, the scalar curvatures are related by:
    $R_{\tilde{g}}=\phi^{-4}R_{\bar{g}}-8\phi^{-5}\Delta_{\bar{g}}\phi=\phi^{-5}(R_{\bar{g}}\phi-8\Delta_{\bar{g}}\phi).$
  2. (ii)
    AM-Lichnerowicz rearrangement: From (40):
    $8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=\Lambda_{J}\phi^{-7}\quad\Rightarrow\quad R_{\tilde{g}}=\phi^{-5}\cdot\Lambda_{J}\phi^{-7}=\Lambda_{J}\phi^{-12}.$
  3. (iii)
    Positivity: Since $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}\geq 0$ and $\phi>0$, we have $R_{\tilde{g}}\geq 0$ automatically.
  4. (iv)
    Strict positivity: $R_{\tilde{g}}>0$ where $\mathcal{S}_{(g,K)}\neq 0$, i.e., where the data deviates from Kerr geometry.
  5. (v)
    Equality case: For Kerr data, $\Lambda_{J}=0$, so $R_{\tilde{g}}=0$ and the monotonicity integrand vanishes.
@@
Key properties (immediate from the construction):
  1. (a)
    $\Lambda_{J}\geq 0$ everywhere (squared norm);
  2. (b)
    $\Lambda_{J}$ is coordinate-independent (tensor norm);
  3. (c)
    $\Lambda_{J}=O(r^{-4-2\tau})$ for asymptotically flat data with decay $\tau>1/2$;
  4. (d)
    For Kerr slices: $E=E^{\mathrm{Kerr}}$ and $B=B^{\mathrm{Kerr}}$ exactly, so $\Lambda_{J}=0$;
  5. (e)
    $\Lambda_{J}=0$ iff $(M,g,K)$ is a Kerr slice (Theorem G.14).
@@
    Regularity at axis: The axisymmetry condition $\mathcal{L}_{\eta}E^{\mathrm{Kerr}}=\mathcal{L}_{\eta}B^{\mathrm{Kerr}}=0$ constrains the harmonic modes to those compatible with axis regularity (only $m=0$ azimuthal modes for the scalar quantities).
@@
Theorem (Mars [83], Bäckdahl–Valiente Kroon [86]): Let $(M,g,K)$ be asymptotically flat, axisymmetric, vacuum initial data. Then $\mathcal{S}_{(g,K)}=0$ if and only if $(M,g,K)$ is isometric to a spacelike slice of the Kerr spacetime.
@@
    Interior extension boundedness: The reference Kerr tensors $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ obtained via Construction B (inward ODE integration) satisfy:
@@
Proof A: Variational Existence (Unconditional—Primary Proof).
@@
where $\beta<0$ is chosen so that functions in $W^{1,2}_{\beta}$ decay exponentially on the cylindrical end. The Euler–Lagrange equation is precisely the AM-Lichnerowicz equation (44).
@@
Proof B: Sub/Super-Solution Method (Conditional on Lemma E.3).
This alternative proof uses the refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$ to establish $\phi\leq 1$ directly.
@@ 6.1 The p-Harmonic Potential
with boundary conditions:
  • At the horizon: $u_{p}|_{\Sigma}=0$, interpreted as $\lim_{t\to\infty}u_{p}(t,y)=0$ along the cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$ (where $t=-\ln s$ and $s$ is distance to $\Sigma$);
  • At infinity: $u_{p}\to 1$ as $r\to\infty$ in the asymptotically flat end.
@@
Critical set structure. The set of critical points $\mathcal{Z}_{p}:=\{x:\nabla u_{p}(x)=0\}$ requires careful analysis because the classical Sard theorem requires $C^{n}$ regularity for functions on $n$-dimensional manifolds, which $C^{1,\beta}$ regularity does not provide. Instead, we use the following specialized results for $p$-harmonic functions:
  1. 1.
    Hausdorff dimension bound (Heinonen–Kilpeläinen–Martio [30]): For $p$-harmonic functions $u:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$, the critical set satisfies $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq n-2$. In dimension $n=3$, this gives $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 1$.
  2. 2.
    Capacitary potentials have isolated critical points (Manfredi [36]): For the AMO potential $u_{p}$ with Dirichlet boundary conditions on connected components, the strong maximum principle combined with saddle point classification shows that $\mathcal{Z}_{p}$ consists of isolated points.
  3. 3.
    Critical value measure zero (consequence): Since $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 0$ for capacitary potentials, the set of critical values $u_{p}(\mathcal{Z}_{p})$ is at most countable, hence has measure zero in $[0,1]$.
@@
Level set behavior at axis intersection points. The level sets $\Sigma_{t}=\{u=t\}$ intersect the axis $\Gamma$ at isolated points (for generic $t$). Near such a point $p\in\Sigma_{t}\cap\Gamma$:
  1. 1.
    The level set $\Sigma_{t}$ is smooth (by implicit function theorem, since $\nabla u\neq 0$ generically);
  2. 2.
    The surface $\Sigma_{t}$ meets the axis orthogonally (by axisymmetry: $\Sigma_{t}$ is rotationally symmetric about $\Gamma$);
  3. 3.
    The mean curvature $H$ and second fundamental form $h$ are bounded at $p$;
  4. 4.
    The Komar integrand $K(\eta,\nu)$ vanishes at $p$ since $\eta=0$ there, so the axis contributes zero to the angular momentum integral.
MOTS-axis intersection. The outermost MOTS $\Sigma$ intersects the axis at exactly two points (the “poles”) by topological considerations ($\Sigma\cong S^{2}$). At these poles:
  • The stability operator $L_{\Sigma}$ has smooth coefficients extending to the poles;
  • The Dain–Reiris inequality $A\geq 8\pi|J|$ accounts for the axis contribution correctly (the proof in [22] handles axis regularity explicitly).
@@
Step 6: Level sets remain in the vacuum region. By hypothesis (H3) of Theorem 1.2, the data is vacuum in the exterior region—i.e., the region $M_{\text{ext}}:=M\setminus\overline{\text{Int}(\Sigma)}$ outside the outermost MOTS satisfies $\mu=|j|=0$. All level sets $\Sigma_{t}$ for $t\in(0,1)$ lie in this exterior region:
  • At $t=0$, $\Sigma_{0}=\Sigma$ is the outermost MOTS (boundary of $M_{\text{ext}}$).
  • For $t>0$, $\Sigma_{t}$ lies outside $\Sigma$ since $u$ increases outward (toward infinity).
  • The monotonicity of $u$ ensures $\Sigma_{t}\subset M_{\text{ext}}$ for all $t\in(0,1)$.
Therefore, the co-closedness condition $d^{\dagger}\alpha_{J}=0$ (equivalently, $d(\star\alpha_{J})=0$) holds throughout the region $\bigcup_{t\in(0,1)}\Sigma_{t}$ swept by the level sets, ensuring the Stokes’ theorem argument applies. ∎
@@
The conservation of angular momentum (Theorem 6.14) relies on the following logic:
  1. 1.
    Setup: Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ on $(M,g)$.
  2. 2.
    Key identity: Vacuum + axisymmetry $\Rightarrow$ $d^{\dagger}\alpha_{J}=0$ (co-closedness).
  3. 3.
    Hodge duality: $d^{\dagger}\alpha_{J}=0$ $\Leftrightarrow$ $d(\star_{g}\alpha_{J})=0$ in 3D.
  4. 4.
    Stokes: $\int_{\Sigma_{t_{2}}}\star\alpha_{J}-\int_{\Sigma_{t_{1}}}\star\alpha_{J}=\int_{W}d(\star\alpha_{J})=0$.
  5. 5.
    Conclusion: $J(t)=J$ constant along the flow.

### v4

@@ 4.3 Axisymmetric Setting
At the poles, $\rho_{\Sigma}=0$, so $\mathcal{T}=O(s\cdot 0)=0$. The explicit constant follows from $\rho_{\Sigma}\leq\rho_{\max}$. Crucially, this estimate holds uniformly on $\Sigma\times[0,s_{0})$, including near the axis. Although the orbit-space coordinates degenerate at the poles, the geometric factor $\rho^{2}$ in the twist term precisely compensates for any potential coordinate singularities, ensuring the perturbation remains $O(s)$ everywhere. This confirms that the twist perturbation is indeed lower-order and does not disrupt the blow-up asymptotics or the cylindrical end structure required for the subsequent analysis. ∎
@@
Lemma 4.16 (Perturbation Stability).
@@
$\|\Phi(v_{1})-\Phi(v_{2})\|_{W^{2,2}_{\beta}}$ $\leq C_{L}\bigl(\|N[v_{1}]-N[v_{2}]\|_{W^{0,2}_{\beta}}$
$\qquad+\|\mathcal{T}[f_{0}+v_{1}]-\mathcal{T}[f_{0}+v_{2}]\|_{W^{0,2}_{\beta}}\bigr).$
@@
Remark 4.17 (Explicit Constants).
The perturbation stability argument involves constants that are explicitly computable from the initial data. In particular, $C_{L}=\|L_{0}^{-1}\|_{W^{0,2}_{\beta}\to W^{2,2}_{\beta}}$ is inversely proportional to the spectral gap $\lambda_{1}(L_{\Sigma})$, so more stable MOTS yield better perturbation control. For axisymmetric vacuum data with strictly stable MOTS, all constants are finite and the argument becomes quantitatively stronger when the MOTS is more stable and the twist is weaker.
@@
$L_{\mathrm{axi}}v$ $=\frac{1}{\sqrt{1+|\nabla f|^{2}}}\biggl[\Delta v-\frac{\nabla^{i}f\nabla^{j}f}{1+|\nabla f|^{2}}\nabla_{ij}v$
$\qquad-(|A_{\Gamma}|^{2}+\mathrm{Ric}(\nu_{\Gamma},\nu_{\Gamma}))v\biggr]+\text{($K$-terms)}+D\mathcal{T}|_{f}\cdot v,$
@@
$0$ $=\int_{\mathcal{C}_{T}}v\cdot L_{\mathrm{axi}}v\,dV_{\bar{g}}$
$=\int_{\mathcal{C}_{T}}\Bigl[-|\nabla v|^{2}+\mathcal{V}v^{2}+O(e^{-\beta_{0}t})|v|^{2}$
$\qquad\qquad+O(e^{-t})|v||\nabla v|\Bigr]dV_{\bar{g}}+\text{(boundary terms)}.$ (22)
@@
  • The blow-up asymptotics $|\nabla f|\sim C_{0}/s$ (see Han–Khuri [29]);
@@
  • The $\rho^{2}$ scaling of the twist term: $\mathcal{T}\propto\rho^{2}$, which vanishes at the poles where $\rho=0$ (Lemmas 4.6 and 4.9).
@@ 5.1 The Conformal Equation
The Jang manifold $(\bar{M},\bar{g})$ has two distinct asymptotic regions requiring different function space frameworks. The asymptotically flat end uses weighted Hölder spaces $C^{k,\beta}_{-\tau}$ with polynomial weight $r^{-\tau}$ (Definition 4.1), while the cylindrical end uses weighted Sobolev spaces $W^{k,p}_{\beta}$ with exponential weight $e^{\beta t}$ (Definition 5.1). These frameworks are compatible on the transition region $\{R_{0}\leq r\leq 2R_{0}\}$ (equivalently $\{0\leq t\leq T_{0}\}$): by Sobolev embedding, $W^{k+1,2}_{\beta}\hookrightarrow C^{k,\beta}$ locally, and both norms are equivalent (up to constants depending on $R_{0}$) on the compact overlap region. This allows elliptic estimates to be “glued” across the transition using standard partition-of-unity arguments. The key point is that the Fredholm index is determined by the asymptotic behavior at both ends, not the transition region.
@@
The modified Lichnerowicz equation with angular momentum terms is:
@@
We verify the sign conventions in the AM-Lichnerowicz equation. Under $\tilde{g}=\phi^{4}\bar{g}$, the scalar curvatures are related by $R_{\tilde{g}}=\phi^{-4}R_{\bar{g}}-8\phi^{-5}\Delta_{\bar{g}}\phi=\phi^{-5}(R_{\bar{g}}\phi-8\Delta_{\bar{g}}\phi)$. Rearranging (24), we have $8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=\Lambda_{J}\phi^{-7}$, which yields $R_{\tilde{g}}=\phi^{-5}\cdot\Lambda_{J}\phi^{-7}=\Lambda_{J}\phi^{-12}$. Since $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}\geq 0$ and $\phi>0$, we have $R_{\tilde{g}}\geq 0$ automatically, with strict positivity $R_{\tilde{g}}>0$ where $\mathcal{S}_{(g,K)}\neq 0$, i.e., where the data deviates from Kerr geometry. For Kerr data, $\Lambda_{J}=0$, so $R_{\tilde{g}}=0$ and the monotonicity integrand vanishes. The existence and uniqueness of the solution $\phi$ in the weighted Sobolev spaces follows from the standard theory for Lichnerowicz-type equations with non-negative source terms, as the negative sign in the operator $L_{AM}$ ensures coercivity (see Theorem 5.10).
@@
Remark 5.5 (Solvability and Sign of Source Term).
The existence and uniqueness of a positive solution $\phi$ to (24) are guaranteed by the structure of the equation. The linear operator $8\Delta_{\bar{g}}+R_{\bar{g}}$ is invertible (Fredholm index 0 and trivial kernel) in the weighted Sobolev spaces $W^{k,p}_{\beta}$ due to the non-negative scalar curvature $R_{\bar{g}}$ and the spectral gap of the MOTS. The nonlinear term $\Lambda_{J}\phi^{-7}$ has the correct sign (negative) to allow for the construction of sub- and super-solutions (e.g., using the maximum principle). Specifically, since $\Lambda_{J}\geq 0$, the term acts as a “sink”, preventing runaway growth. This sign is crucial: a positive source term would lead to potential non-uniqueness or blow-up, but the negative sign ensures the operator is monotone and the solution is unique. The combination of the Fredholm linear theory (Lockhart–McOwen) and the barrier method ensures a unique smooth solution with the prescribed boundary behavior.
@@
Key properties (immediate from the construction): $\Lambda_{J}\geq 0$ everywhere (squared norm); $\Lambda_{J}$ is coordinate-independent (tensor norm); $\Lambda_{J}=O(r^{-4-2\tau})$ for asymptotically flat data with decay $\tau>1/2$; for Kerr slices, $E=E^{\mathrm{Kerr}}$ and $B=B^{\mathrm{Kerr}}$ exactly, so $\Lambda_{J}=0$; and $\Lambda_{J}=0$ if and only if $(M,g,K)$ is a Kerr slice (Theorem G.14).
@@
    Axis regularity: The axisymmetry condition $\mathcal{L}_{\eta}E^{\mathrm{Kerr}}=\mathcal{L}_{\eta}B^{\mathrm{Kerr}}=0$ constrains the harmonic modes to those compatible with axis regularity (only $m=0$ azimuthal modes for the scalar quantities).
@@
Theorem (Kerr Characterization): Let $(M,g,K)$ be asymptotically flat, axisymmetric, vacuum initial data. Then $\mathcal{S}_{(g,K)}=0$ if and only if $(M,g,K)$ is isometric to a spacelike slice of the Kerr spacetime.
@@
Remark 5.8 (Initial Data Characterization).
It is crucial that this characterization applies directly to the initial data $(M,g,K)$ without assuming a priori that the spacetime development is stationary. The condition $\mathcal{S}_{(g,K)}=0$ on the initial slice is sufficient to force the existence of a Killing vector field in the development that makes it stationary (and specifically Kerr).
@@
    Interior boundedness: The reference Kerr tensors $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ obtained via Construction B (inward ODE integration) satisfy:
@@
Proof A: Variational Existence (Unconditional—Primary).
@@
where $\beta<0$ is chosen so that functions in $W^{1,2}_{\beta}$ decay exponentially on the cylindrical end. Specifically, we choose $\beta\in(-\sqrt{\lambda_{1}(L_{\Sigma})},0)$ to ensure the Laplacian is Fredholm (avoiding indicial roots) and to allow the necessary decay for the variational argument. The Euler–Lagrange equation is precisely the AM-Lichnerowicz equation (28).
@@
Proof B: Sub/Super-Solution (Conditional on Lemma E.3). This alternative proof uses the refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$ to establish $\phi\leq 1$.
@@ 6.1 The p-Harmonic Potential
with boundary conditions $u_{p}|_{\Sigma}=0$ (interpreted as $\lim_{t\to\infty}u_{p}(t,y)=0$ along the cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$, where $t=-\ln s$ and $s$ is distance to $\Sigma$) and $u_{p}\to 1$ at infinity in the asymptotically flat end.
@@
Critical set structure. The set of critical points $\mathcal{Z}_{p}:=\{x:\nabla u_{p}(x)=0\}$ requires careful analysis because the classical Sard theorem requires $C^{n}$ regularity for functions on $n$-dimensional manifolds, which $C^{1,\beta}$ regularity does not provide. Instead, we use specialized results for $p$-harmonic functions.
Heinonen–Kilpeläinen–Martio [30] establish that for $p$-harmonic functions $u:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$, the critical set satisfies $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq n-2$, which in dimension $n=3$ gives $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 1$. For the AMO potential $u_{p}$ with Dirichlet boundary conditions on connected components, Manfredi [36] shows that the strong maximum principle combined with saddle point classification implies $\mathcal{Z}_{p}$ consists of isolated points. Consequently, since $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 0$ for capacitary potentials, the set of critical values $u_{p}(\mathcal{Z}_{p})$ is at most countable, hence has measure zero in $[0,1]$.
@@
Level set behavior at axis intersection points. The level sets $\Sigma_{t}=\{u=t\}$ intersect the axis $\Gamma$ at isolated points (for generic $t$). Near such a point $p\in\Sigma_{t}\cap\Gamma$, the level set $\Sigma_{t}$ is smooth (by the implicit function theorem, since $\nabla u\neq 0$ generically), and the surface meets the axis orthogonally (by axisymmetry: $\Sigma_{t}$ is rotationally symmetric about $\Gamma$). The mean curvature $H$ and second fundamental form $h$ are bounded at $p$, and the Komar integrand $K(\eta,\nu)$ vanishes at $p$ since $\eta=0$ there, so the axis contributes zero to the angular momentum integral.
MOTS-axis intersection. The outermost MOTS $\Sigma$ intersects the axis at exactly two points (the “poles”) by topological considerations ($\Sigma\cong S^{2}$). At these poles, the stability operator $L_{\Sigma}$ has smooth coefficients extending to the poles, and the Dain–Reiris inequality $A\geq 8\pi|J|$ accounts for the axis contribution correctly (the proof in [22] handles axis regularity explicitly).
@@
Step 6: Level sets remain in the vacuum region. By hypothesis (H3) of Theorem 1.1, the data is vacuum in the exterior region—i.e., the region $M_{\text{ext}}:=M\setminus\overline{\text{Int}(\Sigma)}$ outside the outermost MOTS satisfies $\mu=|j|=0$. All level sets $\Sigma_{t}$ for $t\in(0,1)$ lie in this exterior region: at $t=0$, $\Sigma_{0}=\Sigma$ is the outermost MOTS (boundary of $M_{\text{ext}}$), and for $t>0$, $\Sigma_{t}$ lies outside $\Sigma$ since $u$ increases outward (toward infinity). The monotonicity of $u$ ensures $\Sigma_{t}\subset M_{\text{ext}}$ for all $t\in(0,1)$. Therefore, the co-closedness condition $d^{\dagger}\alpha_{J}=0$ (equivalently, $d(\star\alpha_{J})=0$) holds throughout the region $\bigcup_{t\in(0,1)}\Sigma_{t}$ swept by the level sets, ensuring the Stokes’ theorem argument applies. ∎
@@
The conservation of angular momentum (Theorem 6.14) relies on the following logic. We start with the Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ on $(M,g)$. The vacuum assumption combined with axisymmetry implies $d^{\dagger}\alpha_{J}=0$ (co-closedness). By Hodge duality in 3D, $d^{\dagger}\alpha_{J}=0$ is equivalent to $d(\star_{g}\alpha_{J})=0$. Applying Stokes’ theorem, we obtain $\int_{\Sigma_{t_{2}}}\star\alpha_{J}-\int_{\Sigma_{t_{1}}}\star\alpha_{J}=\int_{W}d(\star\alpha_{J})=0$, and hence $J(t)=J$ is constant along the flow.